# ANN vs SNN: A case study for Neural Decoding in Implantable Brain-Machine Interfaces

### v1

@@ IV-C Baseline result and effect of filtering
Figure 10: Pareto plots showing trade-offs for different models with increased training data to $80\%$ from the originally used $50\%$: a) Compute cost vs. accuracy b) Memory footprint vs. accuracy
TABLE III: Baseline Performance and Comparison with Prior Works
@@
Sparsity
MACs ACs Memory access
ANN 50% No Filter [17] 0.5818 0.7514 4969.76 0 5,179.46 26.5234
Fwd Filtering 0.5874 0.7514 4971.76 0 5,181.46 26.5312
Block Bid Filtering 0.6064 0.7514 4972.76 0 5,182.46 26.5351
Bid Filtering 0.6168 0.7514 4974.76 0 5,184.46 26.5429
80% No Filter 0.6119 ((((+0.03)))) 0.7417 5000.25 0 5,205.65 26.5234
Fwd Filtering 0.6169 ((((+0.03)))) 0.7417 5002.25 0 5,207.65 26.5312
Block Bid Filtering 0.6361 ((((+0.03)))) 0.7417 5003.25 0 5,208.65 26.5351
Bid Filtering 0.6461 (normal-(((+0.03)normal-))) 0.7417 5005.25 0 5,210.65 26.5429
ANN_3D 50% No Filter [17] 0.6013 0.7348 11507.07 0 11,555.31 134.5234
Fwd Filtering 0.6095 0.7348 11509.07 0 11,557.31 134.5312
Block Bid Filtering 0.6495 0.7348 11510.07 0 11,558.31 134.5351
Bid Filtering 0.656 0.7348 11512.07 0 11,560.31 134.5429
80% No Filter 0.6523 ((((+0.05)))) 0.7324 11676.22 0 11,644.82 134.5234
Fwd Filtering 0.6553 ((((+0.05)))) 0.7324 11678.22 0 11,646.82 134.5312
Block Bid Filtering 0.6769 ((((+0.03)))) 0.7324 11679.22 0 11,647.82 134.5351
Bid Filtering 0.6887 (normal-(((+0.03)normal-))) 0.7324 11681.22 0 11,649.82 134.5429
SNN_3D 50% No Filter [17] 0.6219 0 32256 0 39,057.79 33.1992
Fwd Filtering 0.6318 0 32258 0 39,058.79 33.207
Block Bid Filtering 0.6579 0 32259 0 39,059.79 33.2109
Bid Filtering 0.6687 0 32261 0 39,062.79 33.2187
80% No Filter 0.6564 ((((+0.03)))) 0 32256 0 39,701.38 33.1992
Fwd Filtering 0.6617 ((((+0.03)))) 0 32258 0 39,702.38 33.207
Block Bid Filtering 0.6948 (normal-(((+0.04)normal-))) 0 32259 0 39,703.38 33.2109
Bid Filtering 0.6909 ((((+0.02)))) 0 32261 0 39,706.38 33.2187
SNN_Streaming 50% No Filter 0.6112 0.7453 0 971.26 1,195.28 25.32
Fwd Filtering 0.6109 0.7453 0 973.26 1,197.28 25.3278
Block Bid Filtering 0.6342 0.7453 0 974.26 1,198.28 25.3317
Bid Filtering 0.6458 0.7453 0 976.26 1,200.28 25.3395
80% No Filter 0.6483 ((((+0.04)))) 0.7795 0 883.36 1,044.23 25.32
Fwd Filtering 0.6486 ((((+0.04)))) 0.7795 0 885.36 1,046.23 25.3278
Block Bid Filtering 0.6674 ((((+0.03)))) 0.7795 0 886.36 1,047.23 25.3317
Bid Filtering 0.6761 (normal-(((+0.03)normal-))) 0.7795 0 888.36 1,049.23 25.3395
LSTM 50% No Filter 0.6508 0 22687.97 0 22913.27 90.95
Fwd Filtering 0.6506 0 22690.97 0 22915.27 90.96
Block Bid Filtering 0.6669 0 22690.97 0 22916.27 90.96
Bid Filtering 0.6683 0 22690.97 0 22918.27 90.96
80% No Filter 0.6943 ((((+0.04)))) 0 22687.97 0 22912.40 90.95
Fwd Filtering 0.6932 ((((+0.04)))) 0 22690.97 0 22914.40 90.96
Block Bid Filtering 0.7044 ((((+0.04)))) 0 22690.97 0 22915.40 90.96
Bid Filtering 0.7051 (normal-(((+0.04)normal-))) 0 22690.97 0 22917.40 90.96
SNN 2D[17] 50% - 0.5805 0.9976 0 413.52 - 28.56
rEFH_dynamic [15] $*{}^{*}$ 0.6319 ((((bin_width=128)))) - - - - -
V Discussion
V-A Effect of Reach Removal
As mentioned in Section III-F, some of the reaches in the dataset spanned a much longer duration (sometimes longer than 200 seconds) than the rest which mostly were less than 4 seconds. These reaches (longer than 8 seconds) were removed from training since the NHP was likely unattentive in these cases. However, they were not removed from the testing data and hence, we explored how much improvement in performance is obtained by better curating the test dataset. These results are presented in the Table IV and we can observe that the $R^{2}$ increases by $\approx 0.01$ with the baseline 50% split–the improvement can be much more if other files from [15] are selected. This underlines the effectiveness and necessity of careful data selection from the recordings in [15] while training and testing models.
TABLE IV: Effect of Reach Removal
@@ V-A Effect of Reach Removal
ANN 50% No Filter 0.5921 ((((+0.01))))
Fwd Filtering 0.5977 ((((+0.01))))
Block Bid Filtering 0.6155 ((((+0.009))))
Bid Filtering 0.6249 (normal-(((+0.008)normal-)))
ANN_3D 50% No Filter 0.6133 ((((+0.01))))
Fwd Filtering 0.6225 ((((+0.01))))
Block Bid Filtering 0.6588 ((((+0.009))))
Bid Filtering 0.6690 (normal-(((+0.01)normal-)))
SNN_3D 50% No Filter 0.6286 ((((+0.007))))
Fwd Filtering 0.6437 ((((+0.01))))
Block Bid Filtering 0.6652 ((((+0.007))))
Bid Filtering 0.6755 (normal-(((+0.007)normal-)))
SNN_Streaming 50% No Filter 0.6212 ((((+0.01))))
Fwd Filtering 0.6208 ((((+0.01))))
Block Bid Filtering 0.6433 ((((+0.009))))
Bid Filtering 0.6542 (normal-(((+0.008)normal-)))
LSTM 50% No Filter 0.6784 ((((+0.03))))
Fwd Filtering 0.6779 ((((+0.03))))
Block Bid Filtering 0.6788 ((((+0.01))))
Bid Filtering 0.6800 (normal-(((+0.01)normal-)))
V-B Latency of Filters
Latency between input and output is important for real-time applications with closed-loop operation such as motor decoding. The total time, $T_{tot}$, taken to produce an output by a NN decoder is given by $T_{tot}=St+T_{comp}$ where $St$ is the stride to capture the new input data ($=T_{GT}=4$ ms in this work) and $T_{comp}$ is the time taken to process the computations in the neural network. Given the very fast and energy-efficient In-memory computing (IMC) approaches to implement NN models prevalent now[32, 33] and the small networks considered in this paper, we can assume $T_{comp}<<St$ making the throughput almost entirely dependent on $St$, i.e. time taken to capture new neural input spikes. Note that the bin window, $T_{W}$ does not add any extra penalty on latency of output generation; however, after every change of target, the prediction will be inaccurate for a time related to $T_{W}$ to allow enough relevant input to fill up the bin window. However, output filtering may induce an extra penalty on the latency. Bid filters produced best results as seen in the earlier section; however, they cannot be employed in real-time applications since they need to store the raw data in memory first and then apply forward filtering two times in opposite directions. The block Bid filter is chosen as a compromise where the filter window is used to determine the length or block of samples that are filtered at one time, and the predicted point is located at the center of the sample window. Thus, the latency introduced by the block Bid filter is theoretically equal to half the length of the filter window. In this paper, window size was fixed to 16 samples resulting in a latency of $32$ ms. While this is acceptable in motor decoding, other applications may require lower latency. One possibility is to use analog Bessel filters which do not require such Bid filtering and this will be a focus of future work.
V-C Future Directions
@@
VI Conclusion
Scaling iBMI systems to tens of thousands of channels in the future as well as removing the connecting wires would require significant compression of data on the device to reduce wireless datarates. Integrating the signal processing chain up to the neural decoder offers interesting opportunities to maximize compression. In this context, this work explores the usage of different neural network models and combines them with traditional signal filtering techniques to explore accuracy vs cost trade-offs where the cost is measured in terms of memory footprint and number of operations. Adding Bessel filtering improves the performance of all five NN models with Bidirectional (Bid) and block Bid filtering generating the state-of-the-art results in offline and online filtering respectively. In general, LSTM and SNN_streaming models occupy the high and low ends of the pareto curves (for accuracy vs. memory/operations) respectively while SNN_3D and ANN_3D occupy intermediate positions.
@@ acknowledgment
We acknowledge useful discussions with the Motor decoding group in Neurobench.

### v2

@@ 4.4 Baseline result and Pareto plots
Figure 14: Baseline Pareto plots using $50\%$ training data showing trade-offs for different models: a) Compute cost vs. accuracy b) Memory footprint vs. accuracy. The following colour scheme is used: (1) blue markers are base models without filtering (correspond to results from prior work in [29] for ANN, ANN_3D and SNN_3D) (2) orange markers are models using block Bid filtering with 4th order and block size 32. (3) markers with dark border are tiny variants of base models.
Different ANN and SNN models (baseline and tiny as explained in Section 3) are trained and evaluated on the dataset; the results for the baseline variants are detailed in Table 2. The high accuracy configuration of filter is used here, i.e. the filter order and cutoff frequency are $4$ and $0.05$ for bidirectional filtering, while in block bid filtering, the block size is selected as $32$. Results for the low latency configuration are shown in the Supplementary data.
Computes and model size are obtained from the Neurobench code harness[29] [60]. In NeuroBench, computes are broken down into the following three types: dense, effective Multiply-Accumulates (Effective MACs), and effective Accumulate Synaptic operations (Effective ACs). Dense computes accounts for all zero and nonzero neuronal activations and synaptic connections. This is used to reflect the number of operations needed on hardwares that does not support sparse operations. Effective MACs and effective ACs only take into consideration of operations that are nonzero, i.e. any zero activation by ReLUs, no spike output by SNNs, or zero synaptic connections are ignored, reflecting the operations that would take place on hardwares that support sparsity. NeuroBench computes a model’s footprint by taking the following into consideration: quantization level of the weights, parameters (Weights and Biases), and buffers needed for preprocessing of input data (for a realistic inference comparison). Note for a model’s footprint, zero weights are included as well, as they are part of the connection sparsity metrics.
For a fair comparison across models, we have also added filtering to the output of other models such as ANN and LSTM. The general trend observable from Table 2 and Figure 13 is that filtering improves the $R^{2}$ for all the NN models. Compared to the NN models, signal processing methods like SSKF have much lower computes but the $R^{2}$ is significantly lower due to its inability to track changes in data by varying KF gain[61, 62]. On the other hand, UKF or rEFH have much higher computations due to matrix inversions[30], but still do not attain similar $R^{2}$ as the NN methods. The baseline accuracy we obtained was 0.607 for SNN2 [30], which can be improved to 0.6354 with block filtering. Compared to earlier LSTMs[30, 44] and SNNs[30], the filtered LSTM and SNN_Streaming in this work achieves higher $R^{2}$ with less computes and memory. To visually compare the tradeoffs between accuracy and resource usage, we use pareto plots (Figure 14). The black line is the Pareto frontier, which indicates the best trade-off between the accuracy and operations/memory, the ideal place being the lower right corner of the plot. To keep the plots less cluttered, we only plot results with block Bid filtering since Bid filtering cannot be used in real-time implementations in any case. We also add the tiny variants in the plots. For the pareto plots shown in this section, the following colour scheme is used:
  • orange markers are models using block Bid filtering with block size of 32 and order of 4.
  • markers with dark border are tiny variants of base models
First, we compare the performance using the $50\%$ data split as done in [29] using pareto plots as shown in Figure 14. In terms of the models that forms the pareto front of operations vs. accuracy (Figure 14(a)), we observe that filtered SNNs dominate with SNN_3D occupying the higher part while SNN_Streaming occupying the lower part. These results can be taken as a gold standard for the neurobench suite[29] at this time since they represent the highest reported accuracy so far. The two SNNs variants show a big difference in terms of operations required ($\approx 100$x) and accuracies ($\approx 4\%$). The block filtering increased the accuracy of SNN_3D and SNN_Streaming by 8% and 5.7% with only a 0.015% and 0.512% increase in computes, respectively. Figure 15 plots the actual trajectory of a ground truth reach waveform, a prediction each from SNN_3D and SNN_Streaming, and the corresponding filtered versions. It can be seen how the filtered waveform is smoother and resembles the more natural motion of the primate’s finger. Also. outputs from SNN_Streaming are inherently smoother than SNN_3D due to its internal memory through membrane potentials which do not get reset like SNN_3D. In terms of memory usage (Figure 14(b)), the pareto front is also dominated by filtered SNNs. Here, block filtering usage only increases model size by a mere 0.058% and 0.077% for SNN_3D and SNN_Streaming. The ANN_3D models have highest memory usage due to their input dimension being expanded by $m$ times to $m\times N_{ch}$–the weights in the first layer are dominant for memory footprint since $N_{0}>>N_{1},N_{2},N_{3}$.
Looking deeper at the effects of filtering, we see that SNN_Streaming with block Bid filtering achieves similar accuracy of $\approx 0.64$ as the LSTM without filtering at $\approx 27\%$ memory and $\approx 25\%$ computations. This confirms our initial hypothesis that adding memory via filtering to SNN models can indeed make their performance similar to recurrent ANNs. Even the tiny variant of SNN_3D model achieves higher accuracy with slightly less operations and only $30\%$ memory usage compared to LSTM. In summary, filtered SNNs are the best performing models and either SNN_3D or SNN_Streaming may be chosen depending on the desired tradeoff between accuracy and resource usage.
Figure 15: Predicted trajectory of X-velocity with and without filter showing the smoothness introduced by the filter making it similar to natural motion: a) SNN_3D. b) SNN_Streaming
4.5 80% vs 50% Training Split
To assess the performance of models when the training data increases, we increase the baseline training data from 50% to 80% as done in [40]. The results are listed in Table 2 and plotted in Figure 16. As expected, the $R^{2}$ of all models is generally higher by $0.03-0.04$ (written in parenthesis in Table 2) compared to the 50% baseline training data, which shows a high capacity for future improvement with more data. Similar to Figure 14, both the pareto curves in Figure 16 for computations and memory are dominated by SNNs with block Bid filtering. However, one filtered LSTM model and one filtered ANN_3D model are now placed on the pareto frontier for computations.
Figure 16: Pareto plots showing trade-offs for different models with increased training data to $80\%$ from the originally used $50\%$: a) Compute cost vs. accuracy b) Memory footprint vs. accuracy. The following colour scheme is used: (1) blue markers are base models without filtering (2) orange markers are models using block Bid filtering with 4th order and block size 32. (3) markers with dark border are tiny variants of base models.
5 Discussion
5.1 Effect of Reach Removal
As mentioned in Section 3.6, some of the reaches in the dataset spanned a much longer duration (sometimes longer than 200 seconds) than the rest which mostly were less than 4 seconds. These reaches (longer than 8 seconds) were removed from training since the NHP was likely unattentive in these cases. However, they were not removed from the testing data and hence, we explored how much improvement in performance is obtained by better curating the test dataset. These results are presented in the Table 4 and we can observe that the $R^{2}$ increases by $\approx 0.01$ with the baseline 50% split–the improvement can be much more if other files from [19] are selected. This underlines the effectiveness and necessity of careful data selection from the recordings in [19] while training and testing models.
Table 4: Effect of Reach Removal–increase in $R^{2}$ over baseline for high accuracy filter configuration shown in parenthesis
Models Split Filters R2
ANN 50% No Filter 0.5921 ((((+0.01))))
ANN_3D 50% No Filter 0.6133 ((((+0.01))))
SNN_3D 50% No Filter 0.6286 ((((+0.007))))
Block Bid Filtering 0.6784 ((((+0.006))))
Bid Filtering 0.6755 ((((+0.007))))
SNN_Streaming 50% No Filter 0.6212 ((((+0.01))))
Block Bid Filtering 0.6532 ((((+0.008))))
Bid Filtering 0.6542 ((((+0.008))))
LSTM 50% No Filter 0.6784 ((((+0.03))))
5.2 Generalization to other datasets
To show the general applicability of our work, we have evaluated the proposed methods on another neural decoding dataset referred to as ‘MC Maze’[63, 64] that has been used to evaluate other SNN decoders [45, 65, 66]. Briefly, this dataset contains recordings from the motor and premotor cortex of a monkey as it performed delayed reaching tasks [64]. The reaches were either straight or curved to avoid virtual barriers. Neural data was recorded using two 96-electrode arrays implanted in the PMd and M1 regions. After an offline spike sorting, spike information of 107 neurons, hand position, and monkey gaze position in 1 ms bins are provided. Reaching tasks last up to 600 ms.
Figure 17: Comparison of results for ANN, ANN_3D, SNN_3D, SNN_Streaming and LSTM with CL_SNN and Kalman filter on the MC Maze dataset. Except the LSTM, all other models show statistically significant improvement due to filtering.
All the models described in the earlier sections were trained on this task. Decoders were trained on spiking data from 130 ms prior to movement onset to 370 ms after movement onset[64]. A block bidirectional filtering with a block size of 32 is implemented. Figure 17 displays the results of a comparison with the offline decoding results from [45] using a Kalman Filter and a continuous learning SNN (CL_SNN). To align with the results reported in [45], training data is set as 60% while others are halved and set to testing/validation. It can be seen that all the decoding methods described here performed better than the earlier work[45] (note that the X and Y velocity regressions are not shown separately as in [45], the final results are represented as the average of the X and Y directions (i.e. $R^{2}=\frac{R_{X}^{2}+R_{Y}^{2}}{2}$), with the calculation details provided in the Section 3.5). Moreover, filtering improves decoding accuracy in a statistically significant way (paired t-test) for all the models except LSTMs. SNN_3D again achieves the best performance after filtering while SNN_Streaming has the best tradeoff.
5.3 New training methods
The results in the earlier works were obtained by training the models directly for regression using backpropagation. However, regression is generally considered a more challenging task compared to classification. Hence, there is potential for exploring other training techniques such as Cascade Classification Based Regresion (CCBR), that cast the regression problem into a framework for classification [67]. We describe the method briefly here, with details in [67]. The output space is divided into zones (different classes) and the first classifier predicts the sample should fall in which zone. This produces a regression result corresponding to the centroid of the zone. Following this, the next set of classifiers predicts the remaining error in regression. We did some preliminary investigation and tested the CCBR method in the same way as the previous models, and the result is shown in Table 5, where the classifier is chosen as ANNs/SNNs instead of an SVM as in [67]. In most cases, the accuracy of CCBR is lower than that of the original model. Nevertheless, we still see the same two trends: (1) Block bid filtering can significantly improve accuracy. (2) SNN_3D with filtering achieves best accuracy while SNN_Streaming achieves best tradeoff between accuracy and resource usage.
Table 5: The performance of CCBR training.
Activation Model Size
Models Split Filters R2 Sparsity Computes (((( kB ))))
MACs ACs Memory
Access
@@ 5.3 New training methods
Block Bid Filtering 0.5716 0.7514 11429 0 9315.86 95.72
@@
Block Bid Filtering 0.5995 0.7348 55781 0 19483.228 442.23
[…]
Block Bid Filtering 0.5094 0.7453 0 1947 2333.3 91
[…]
Block Bid Filtering 0.6044 0 51749 0 65341.4 116.27
5.4 Future Directions
@@
The results presented in this work were based on offline decoding while in real life, experiments are performed in a closed-loop mode with visual feedback being commonly used[68, 69]. A software tool, online prosthetic simulator (OPS)[69] has been developed to emulate this closed-loop operation in real experiments and we will use this as our next step of exploration. Furthermore, the characteristics of the acquired signal also change over time due to scar tissue formation on electrodes or micro-motion of electrodes[70] We will also investigate the possibility of using continuous learning [45] to address the issue of data drift.
@@
6 Conclusion
Scaling iBMI systems to tens of thousands of channels in the future as well as removing the connecting wires would require significant compression of data on the device to reduce wireless datarates. Integrating the signal processing chain up to the neural decoder offers interesting opportunities to maximize compression. In this context, this work explores combining SNNs with traditional signal filtering techniques to improve their accuracy vs cost trade-offs where the cost is measured in terms of memory footprint and number of operations. Adding Bessel filtering improves the performance of both types of SNN models and block Bidirectional filtering generating the state-of-the-art results. Two filter variants for high accuracy and low latency are shown. In general, filtered SNN_3D and filtered SNN_Streaming models occupy the high and low ends of the pareto curves (for accuracy vs. memory/operations) respectively. Filtering the output of both ANN and SNN models with the high accuracy variant of Bessel filters exhibited statistically significant improvement in accuracy.
@@ Acknowledgment
We acknowledge useful discussions with the Motor decoding group in Neurobench. The work done in this paper was partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. CityU 11200922).
7 Supplementary Data
.1 Pareto plots for low-latency filter configuration with order 2 and block size 16
Figure A.18: Pareto plots with 50% data split for low--latency Bessel filter configuration with block size of 16 and order of 2. (a) Compute cost vs. accuracy (b) Memory footprint vs. accuracy. SNNs with filtering dominate both pareto plots with one LSTM model appearing in the computation pareto. The following colour scheme is used: (1) blue markers are base models without filtering (2) orange markers are models using block Bid filtering with 2nd order and block size 16. (3) markers with dark border are tiny variants of base models.
Figure A.19: Pareto plots with 80% data split for low--latency Bessel filter configuration with block size of 16 and order of 2. (a) Compute cost vs. accuracy (b) Memory footprint vs. accuracy. SNNs with filtering dominate both pareto plots with one LSTM model appearing in the computation pareto. The following colour scheme is used: (1) blue markers are base models without filtering (2) orange markers are models using block Bid filtering with 2nd order and block size 16. (3) markers with dark border are tiny variants of base models.
.2 Effect of 3 filters on 5 sample reaches from 5 different files
Figure A.20: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$indy\_20160630\_01"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.
Figure A.21: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$indy\_20170131\_02"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.
Figure A.22: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$loco\_20170210\_03"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.
Figure A.23: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$loco\_20170215\_02"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.
Figure A.24: Comparison of three types of filters in terms of velocity and position based on one reach in the file "$loco\_20170301\_05"$. The three columns indicate SNN with Bessel, Butter, and Chebyshev filter respectively. The first row shows the X velocity, the second row shows the Y velocity. 4-th order filter with block size of 32 and cut-off frequency of 0.05 is used.